OpenStudy (anonymous):

Please The length of a rectangle is three feet shorter than twice its width. If the area of the rectangle is 170ft^2, is what is the length of the rectangle?

OpenStudy (allank):

Let the width be w. thus the length is (2w-3). thus 170=w(2w-3) 170=2w^2-3w 0=2w^2-3w-170 Solve that quadratic to get the width w. Then calculate the length. I hope that helps.

OpenStudy (anonymous):

ok.... gimme a sec... solving

OpenStudy (allank):

I have to logout, so I hope you got w=10.

OpenStudy (anonymous):

yes.... thank you... solve by completing the square?

OpenStudy (allank):

You're welcome. Yes, completing the square will work, but is quite a long process. I prefer using the quadratic formula.

OpenStudy (anonymous):

really quick... whats the formula

OpenStudy (anonymous):

completing the square i kept getting wrong answers

OpenStudy (allank):

I recommend reading about both completing the square and the quadratic formula (which is essentially a short-cut of completing the square) here: http://tutorial.math.lamar.edu/Classes/Alg/SolveQuadraticEqnsII.aspx

OpenStudy (anonymous):

thank you

OpenStudy (allank):

You're welcome. Do you still want the quadratic formula?

OpenStudy (allank):

\[x=(-b +/- \sqrt{b^2-4ac})/2a \]

OpenStudy (allank):

That hasn't come out quite as it's supposed to, but you'll get the exact one in the link I've posted. Best wishes.